Follow the process of completing the square to solve 2x^2 + 8x - 12 = 0.
How will the left side of the equation factor in step 5?
(2x + 32)^2
(4x + 8)^2
(4x + 16)^2

Answer by Alan3354(69443)

Follow the process of completing the square to solve 2x^2 + 8x - 12 = 0.
How will the left side of the equation factor in step 5?
-----
2x^2 + 8x - 12 = 0
x^2 + 4x - 6 = 0
x^2 + 4x = 6
x^2 + 4x + 4 = 6 + 4
x^2 + 4x + 4 = 10
(x+2)^2 = 10